3<=|x-2|<=5
my answer which is marked wrong was [5,7]
should it be {5,6,7} (-infinity,7] n [5,infinity) ?

Found 2 solutions by rfer, edjones:
Answer by rfer(16322)

It looks like your answer should be between 3 & 5

Answer by edjones(8007)

3<=|x-2|
.
3<=x-2
5<=x
.
-3>=x-2
-1>=x
.
|x-2|<=5
.
x-2<=5
x<=7
.
x-2>=-5
x>=-3
.
So:
-3<=x<=-1, 5<=x<=7
or
[-3, -1], [5, 7]
.
Ed